Pursley, Inc. owns 70 percent of Harry Corp. The consolidated income statement for a year reports $50,000 Noncontrolling Interest in Harry Corp.'s Net Income. Harry paid dividends in the amount of $80,000 for the year. What are the effects of these transactions in the consolidated statement of cash flows for the year?

Financing Activities Operating Activities
A) Increased by $24,000 Increased by $15,000
B) Decreased by $15,000 Unaffected
C) Unaffected Decreased by $15,000
D) Decreased by $24,000 Unaffected
E) Unaffected Increased by $24,000
Answer #1

Answer:

D) Decreased by $24,000 Unaffected

The dividend paid would get reduced to the extent of the controlling interest of 80000*70% = $56000. So cash outflow = 80000-56000 = $24,000 under financing activities.